nolegirl24:

A polynomial function can be written as (x + 2)(x + 3)(x − 5). What are the x-intercepts of the graph of this function? (2, 0), (3, 0), (−5, 0) (−2, 0), (−3, 0), (5, 0) (2, 0), (3, 0), (5, 0) (−2, 0), (−3, 0), (−5, 0)

nolegirl24:

@ThisGirlPretty help :(

nolegirl24:

I think i got it but i wanna make sure

dude:

Set each factor in the parenthesis equal to 0 x + 2=0 x + 3=0 and x − 5=0

nolegirl24:

Yes!!! So I am on the right track :O

dude:

Yay :P

nolegirl24:

So it would be (-2,0)(-3,0)(5,0)?

dude:

Right

nolegirl24:

YAY! I got it :O

dude:

Noice ;p